# Using quantum nonlocality for device-independent confirmation of relativistic effects

### v1

@@ Abstract
Synchronizing clocks to measure time is a fundamental process underpinning every practical communication task from GPS to parallel computation. However, as the current protocols are based on classical communication between the sender and receiver, they are prone to simple attacks that could cause a slight delay in the signal which would then cause a massive error in further operations. In this work, we first explain a simple attack that in principle can cause an arbitrary delay in the signal between sender and receiver. We then propose a way to overcome this problem by using a recently contrived idea of device-independent certification which utilises quantum nonlocality. Consequently, clocks can be synchronized in a highly secure way without trusting any devices in the setup. We then further extend this proposal to observe relativistic time dilation in a device-independent manner.
@@
Synchronizing clocks— Before proceeding, let us clarify that synchronizing clocks in this work implies synchronizing the time interval of some event among spatially separated parties and not the absolute time of the clocks. For simplicity, we will assume that the absolute time of the clocks is the same at the start of the protocol. A simple way to do this is to bring the two clocks to the same location and then put the same time on both clocks. Moreover, we assume in this section that the clocks are rest with respect to each other. Here we focus on whether the time interval of some event measured by these clocks can be manipulated by some attacker which will also cause an error in recording their position with respect to each other. For instance, if the time of two clocks are denoted by $t_{1},t_{2}$, then we want to ensure that if any of the two clocks record a time interval of some event as $\Delta t_{i}(i=1,2)$ in their frame of reference, then we want to ensure $\Delta t_{1}=\Delta t_{2}$ in a device-independent way, that is, without trusting or knowing the internal workings of any device involved in the experiment.

### v2

@@ Abstract
Synchronizing clocks to measure time is a fundamental process underpinning every practical communication task from GPS to parallel computation. However, as the current protocols are based on classical communication between the sender and receiver, they are prone to simple attacks that could cause a slight delay in the signal which would then cause a massive error in further operations. In this work, we first explain a simple attack that in principle can cause an arbitrary delay in the signal between sender and receiver. We then propose a way to overcome this problem by using a recently contrived idea of device-independent certification which utilises quantum nonlocality. Consequently, clocks can be synchronized in a highly secure way without trusting any devices in the setup. We then extend this proposal to observe relativistic time dilation in a device-independent manner.
@@
Synchronizing clocks— Before proceeding, let us clarify that synchronizing clocks in this work implies synchronizing the time interval of some event among spatially separated parties and not the absolute time of the clocks. For simplicity, we will assume that the absolute time of the clocks is the same at the start of the protocol. A simple way to do this is to bring the two clocks to the same location and then put the same time on both clocks. Moreover, we assume in this section that the clocks are rest with respect to each other. Here we focus on whether the time interval of some event measured by these clocks can be manipulated by some attacker which will also cause an error in recording their position with respect to each other. For instance, if the time of two clocks are denoted by $t_{1},t_{2}$, then if any of the two clocks record a time interval of some event as $\Delta t_{i}(i=1,2)$ in their frame of reference, then we want to ensure $\Delta t_{1}=\Delta t_{2}$ in a device-independent way, that is, without trusting or knowing the internal workings of any device involved in the experiment.
@@
Some implicit assumptions in the above scenario—
  • (Assumption 1:) Eve has no quantum memory; thus, she can not store the quantum bit sent by the source and release it at a later time.
  • (Assumption 2:) Authenticated classical channel is required (like quantum cryptography) to ensure that the commnication is genuinely between a satellite and the user, and not between an adversary and the user. This stops impersonation attacks by the adversary.
Given these two assumptions, one can ensure that there are no time-delay attacks in the GPS using the above scenario. Both of these assumptions might be well-justified in some particular scenario, but the scheme is not device-independent unless assumption 1 can be removed.